OpenStudy (anonymous):

a^2+b^2=10^2 What will work for this, what would a and b be?

OpenStudy (blockcolder):

5sqrt(2) for a and b would work.

OpenStudy (anonymous):

but that would be 50 instead of 100

OpenStudy (blockcolder):

\[(5\sqrt{2})^2=50\\ 50+50=100\]

OpenStudy (paxpolaris):

the only integral solutions ... i think are: a=6, b=8 a=8, b=6

OpenStudy (anonymous):

I was looking for a simpler answer but thank you

OpenStudy (anonymous):

directed o blockcolder

OpenStudy (anonymous):

*to

Directrix (directrix):

@smoothiebit123 I'm thinking there may be infinitely many points that work, all located on a circle of radius 10.

Directrix (directrix):

Center of circle would be (0,0). Question: Do you want only integer values of a and b?

OpenStudy (anonymous):

i was doing Pythagorean therom

Directrix (directrix):

Regardless, there are infinitely many solutions that are positive a and the corresponding positve b which lie on the quarter circle in quadrant one of the circle. The Pythagorean Theorem, the formula for a circle, and the distance formula are inextricably related. Bottom Line: There is no unique a and b which are solutions to the equation you posted.